Fermi National Laboratory


Accelerator Update

Friday November 19

The shutdown ends on Wednesday, December 1.

Tevatron

Completed Work
Installed Deep rod monuments at B0, C0 and D0
HLS system at D0 low beta quads
Recheck A-48 collimator
Replaced kicker at F17
Repaired voltage leads on D-49 spool
Cable pulls for TEL 2
Clean up of C0, D0, and A0
General inspection

Linac

The Linac is up and running.

Booster

Booster is up and running.

Antiproton Source

Debuncher Injection Region - Vacuum pump down.

AP1/3 Line -Waiting on D:H926 testing.

AP2 Line - Beam stop plan: Replace old beam stop cavity with steel shielding around evacuated beam pipe centered on AP2 beam line. Install new beam stop in accessible portion of AP2 line.

SO Cord Remediation - Complete.

D/A Lines - Vacuum good. The new septum has been realigned.

Main Injector/Recycler

Some quadrupoles need realignment.

The Main Injector work is done except for MI-30.

Electron Cooling Shutdown Work - All components installed

3 of 4 beam lines baked and under final vacuum
4th beam line (Cooling Section) leak tight and ready for bake
Recycler vacuum is common to Cooling Sections and has leaks
Leaks may be repaired by Friday morning - 11/19/04
If repaired, may start bake today - 11/19/04

If bake starts today, we could be out by 11/24, if all goes well

~12 hours needed after bake to finish up and vacate tunnel

MI beam line reinstallation

Should be complete today 11/19/04
Alignment and pump down to follow

MI Bus clamping issue

Retrofit should be complete today 11/19/04
Testing this weekend
Several minor tasks over the next couple of days will be completed

MiniBooNE

Rigging completed. Conducting instrumentation testing. Final alignment schedule for next week. The experiment will be ready for beam by November 30.

NuMI

Final alignment problem due to wall monuments, but work has been completed. Vacuum pump down complete. Power supply testing and polarity checks needed. The experiment will be ready for a December 3rd start up.

Cryogenic System

There's an air to helium leak on the sector A49 spool piece; the options are to run with it, fix it, or to replace the spool piece, but nothing has been decided. There's a vacuum problem at TeV sector E2. There's a fault to ground on the lead for D4H2S. A cold compressor at D4 service building had to be replaced. The TeV cryo system is cold and the TeV has been ramped to 980 GeV.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 11/22/2004   email Fermilab